Physical Chemistry

if an electron is accelerated through 20V potential difference if it has already 5 eV kinetic energy then wavelength of the electron is ?

Divya
7 Years agoGrade

1 Answer

Nikhil
7 Years ago
we know that work done = eV
so final K.E. = 25eV
wavelength(in angestrom) = 12400/k.e.(in eV)  
                                          =  12400/25 = 496 angestrom